OpenStudy (anonymous):

Given f(x)=2x-2, which is an ordered pair, (x,y) for the element in the domain x=-3?

OpenStudy (anonymous):

f(-3)=2*-3-2 f(-3)=-6-2 f(-3)=-8 so y=-8 and x=-3 so the ordered pair is (-3,-8)

OpenStudy (anonymous):

Thanksss

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

Hey can you help me with another problem?